OpenStudy (anonymous):

. Write an equation of a line that passes through the points (2,3) and (2,-6). Hint: To answer this question, first find the slope using these 2 points.

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=-9/0= undefined x-2=0 is the equation

OpenStudy (anonymous):

thanks